# Milky Way-est: Cosmological Zoom-in Simulations with Large Magellanic Cloud and Gaia–Sausage–Enceladus Analogs

### v1

@@ Abstract
We present Milky Way-est, a suite of 20 cosmological cold-dark-matter-only zoom-in simulations of Milky Way (MW)-like host halos. Milky Way-est hosts are selected such that they: ($i$) are consistent with the MW’s measured halo mass and concentration, ($ii$) accrete a Large Magellanic Cloud (LMC)-like ($\approx 10^{11}~{}{M_{\odot}}$) subhalo within the last $1.3~{}{\rm Gyr}$ on a realistic orbit, placing them near $50~{}{\rm kpc}$ from the host center at $z\approx 0$, and ($iii$) undergo a $>$1:5 sub-to-host halo mass ratio merger with a Gaia–Sausage–Enceladus (GSE)-like system at early times ($30.67<z<$ 3). Hosts satisfying these LMC and GSE constraints constitute $<1\%$ of all halos in the MW’s mass range, and their total masses grow rapidly at late times due to LMC analog accretion. Compared to hosts of a similar final halo mass that are not selected to include LMC and GSE analogs, Milky Way-est hosts contain $22\%$ more subhalos with present-day virial masses above $10^{8}~{}{M_{\odot}}$ throughout the virial radius, on average. This enhancement reaches $\approx 80\%$ in the inner $100~{}\mathrm{kpc}$ and is largely, if not entirely, due to LMC-associated subhalos. These systems also induce spatial anisotropy in Milky Way-est subhalo populations, with $\approx 60\%$ of the total subhalo population within $100~{}{\rm kpc}$ found in the current direction of the LMC. Meanwhile, we find that GSE-associated subhalos do not significantly contribute to present-day Milky Way-est subhalo populations. These results provide context for our Galaxy’s dark matter structure and subhalo population and will help interpret a range of measurements that are currently only possible in the MW.
@@ 1 Introduction
The Milky Way (MW)’s dark matter distribution is a cornerstone of near-field cosmology. In recent years, astrometric measurements (Gardner et al., 2021), photometric surveys (Drlica-Wagner et al., 2020, 2022), and spectroscopic campaigns (Majewski et al., 2017; Naidu et al., 2020; Cooper et al., 2023) have qualitatively advanced our understanding of the events that contributed to the buildup of dark matter in the MW. These observations reveal that the MW’s most massive surviving and disrupted satellites—namely, the recently accreted Large Magellanic Cloud system (LMC; van der Marel et al. 2002; Besla et al. 2007; Kallivayalil et al. 2013; Vasiliev 2023) and the ancient Gaia–Sausage–Enceladus merger (GSE; Belokurov et al. 2018; Helmi et al. 2018)—play a crucial role in shaping its dark matter distribution and substructure (e.g., Helmi 2020; Conroy et al. 2021).
@@
In particular, the GSE merger is a key event in the formation of the Milky Way galaxy, as evidenced by the chemo-kinematic structure of the MW’s disk (e.g., Grand et al. 2020; Belokurov & Kravtsov 2022; Semenov et al. 2024; Dillamore et al. 2024). GSE also significantly impacts the dynamical structure of the MW stellar halo, which is dominated by stars on radial orbits deposited by this merger at early times (e.g., Fattahi et al. 2019; Dillamore et al. 2022; Han et al. 2022). Meanwhile, the LMC drives the MW’s dark matter distribution away from dynamical equilibrium, inducing wakes, deformation, and reflex motion that have recently been detected (e.g., Garavito-Camargo et al. 2019, 2021; Petersen & Peñarrubia 2021; Conroy et al. 2021; Lilleengen et al. 2023). The LMC also brings its own substructure into the MW, reflected by the existence of kinematically associated satellite galaxies (e.g., Kallivayalil et al. 2018; Patel et al. 2020; Samuel et al. 2021) and spatial anisotropy in the distribution of MW satellite galaxies (e.g., Jethwa et al. 2016; Nadler et al. 2020).
@@
From a theoretical perspective, cosmological simulations provide detailed predictions for the growth and structure of MW-mass halos, including their subhalo populations (e.g., Springel et al. 2008; Diemand et al. 2008; Mao et al. 2015; Griffen et al. 2016; Sawala et al. 2016; Wetzel et al. 2016; Poole-McKenzie et al. 2020; Nadler et al. 2023b). Most studies that use simulations to interpret MW observations select hosts with masses and concentrations consistent with those inferred for the MW halo, regardless of merger history. In hydrodynamic contexts, selection often includes central galaxy properties (e.g., stellar mass and morphology). Other simulations require MW analogs to reside in environments similar to the Local Group or Local Volume (e.g., Garrison-Kimmel et al. 2014; Carlesi et al. 2016; Sawala et al. 2022).
@@
Motivated by the observational advances discussed above, recent simulations have focused on halos that satisfy additional constraints on the MW’s merger history. Considering GSE, Bignone et al. (2019) and Bose et al. (2020) study hosts in cosmological hydrodynamic simulations that include GSE analogs, Rey et al. (2023) present zoom-in simulations that are engineered to include GSE analogs with controlled properties, and García-Bethencourt et al. (2023) present a high-resolution hydrodynamical simulation that includes analogs of the GSE along with two other early mergers, Kraken and Sequoia. Considering the LMC, several authors have analyzed small numbers of MW-mass hosts with LMC analog subhalos using gravity-only (e.g., Erkal & Belokurov 2020; Sales et al. 2017) or hydrodynamic simulations (e.g., Barry et al. 2023; Smith-Orlik et al. 2023; Arora et al. 2023). Beyond individual constraints, Evans et al. (2020) analyze MW-like hosts with both LMC and GSE analogs using large-volume cosmological simulations, and Nadler et al. (2020) study zoom-ins of two MW-mass hosts with analogs of both systems. Despite this progress, no existing simulation suite is designed to study MW-mass hosts that accrete an LMC analog at late times and undergo a major merger with a GSE analog at early times. Thus, it is timely to develop high-resolution cosmological simulations that incorporate both accretion events.
@@
Here, we introduce Milky Way-est, a suite of $20$ cosmological cold-dark-matter-only zoom-in simulations focused on hosts that are consistent with the MW’s halo mass and concentration measurements and that accrete both LMC and GSE analogs. Given our LMC and GSE analog selection criteria, we find that $<1\%$ of halos in the relevant mass range undergo both accretion events, adding to previous findings regarding the rarity of the Magellanic Clouds (e.g., Boylan-Kolchin et al. 2010; Busha et al. 2011; Liu et al. 2011). To benchmark our results, we compare Milky Way-est to the Symphony Milky Way-mass zoom-in simulation suite (hereafter Symphony MW; Mao et al. 2015; Nadler et al. 2023b). Symphony MW includes $45$ MW-mass hosts, run with the same resolution and numerical settings as Milky Way-est, which are not selected to include LMC and GSE analogs. Comparing Milky Way-est and Symphony MW hosts can, therefore, isolate the effects of our GSE and LMC constraints on hosts’ mass accretion histories (MAHs) and subhalo populations. We will show Milky Way-est hosts’ formation times, subhalo mass functions (SHMFs), and subhalo radial and spatial distributions significantly differ from Symphony MW. These differences are largely driven by our LMC analog systems. Thus, it is important to model both the LMC and GSE to accurately interpret measurements of the MW’s dark matter distribution and substructure.
@@
Milky Way-est simulations resolve subhalos down to present-day masses of $\approx 10^{8}~{}{M_{\odot}}$, near the threshold of galaxy formation (Benitez-Llambay & Frenk, 2020; Nadler et al., 2020); detailed comparisons between Milky Way-est and Symphony MW will therefore help contextualize upcoming observations of the MW satellite population, which probe galaxy formation and dark matter physics (Nadler et al., 2024). More generally, Milky Way-est provides a realistic setting for testing empirical galaxy–halo connection models (e.g., Nadler et al. 2019; Wang et al. 2021), conducting zoom-in simulations of galaxy formation (e.g., Wetzel et al. 2016; Font et al. 2020; Applebaum et al. 2021), and exploring physics beyond the cold dark matter paradigm (e.g., Nadler et al. 2021b, a; Mau et al. 2022; Yang et al. 2023) in a setting that facilitates accurate predictions for near-field cosmology.
@@
This paper is organized as follows. Section 2 describes our simulations, including host halo selection criteria and the Milky Way-est sample. In Section 3, we present the properties of Milky Way-est hosts, LMC analogs, and GSE analogs. In Section 4, we study Milky Way-est host halo MAHs and subhalo populations compared to Symphony MW. We discuss the implications of our results, areas for future work, and caveats in Section 5; we conclude in Section 6. We list the names and properties of Milky Way-est systems (including their LMC and GSE analog halos) in Appendix A, and a convergence test using one higher-resolution Milky Way-est resimulation is presented in Appendix B.
@@ 2 Simulations
We begin by describing our simulations, including technical details (Section 2.1), host halo selection criteria (Section 2.2), and the Milky Way-est sample that we resimulate using the zoom-in technique (Section 2.3).
@@
Following Nadler et al. (2023b), all of our simulations use $h=0.7$, $\Omega_{\rm m}=0.286$, $\Omega_{\Lambda}=0.714$, $\sigma_{8}=0.82$, and $n_{s}=0.96$ (similar to WMAP 9; Hinshaw et al. 2013). Halo catalogs and merger trees are generated using Rockstar and Consistent-Trees (Behroozi et al., 2013b, a). We define “peak” as the highest value of a quantity along a halo’s main branch, and we measure concentration using the Navarro–Frenk–White (NFW; Navarro et al. 1997) definition.
2.1 Overview
@@
Milky Way-est hosts are selected from the Chinchilla c125-1024 cosmological volume based on several criteria, including host halo properties that are consistent with inferred properties of the MW’s dark matter halo and the existence of realistic LMC and GSE analogs, as described in Section 2.2. c125-1024 is a dark matter-only simulation run with a box length of $125~{}{\rm Mpc}~{}h^{-1}$ with $1024^{3}$ particles (see Mao et al. 2015 for details). We generate zoom-in initial conditions with four refinement regions, yielding an equivalent of $8192$ particles per side in the most refined region, using MUSIC (Hahn & Abel, 2011). The highest-resolution region corresponds to the Lagrangian volume of particles within ten times the $z=0$ virial radius of each host ($R_{\mathrm{vir,host}}$) in c125-1024 (with the exception of Halo756, for which we use $7R_{\mathrm{vir,host}}$ to avoid an expensive resimulation). [^1] Virial quantities are calculated using the Bryan & Norman (1998) overdensity, which corresponds to $\Delta_{\mathrm{vir}}\approx 99$ times the critical density of the universe in our cosmology at $z=0$.
@@
The numerical parameters of our zoom-in simulations match those for Symphony MW (see Nadler et al. 2023b for details). The dark matter particle mass in the highest-resolution region is $m_{\mathrm{part}}=4.0\times 10^{5}~{}{M_{\odot}}$ and the comoving Plummer-equivalent gravitational softening is $\epsilon=170~{}{\rm pc}~{}h^{-1}$. We also verify convergence using a resimulation of one Milky Way-est host with 8×8$\times$8 × smaller particle mass and 2×2$\times$2 × smaller softening (see Appendix B). We initialize each zoom-in at $z=99$ and save 236 snapshots until $z=0$. At late times, the typical time spacing between snapshots is $25~{}{\rm Myr}$; as a result, we run some Milky Way-est zoom-ins slightly past $z=0$ to ensure that our LMC analogs reach a distance of $\approx 50~{}{\rm kpc}$ (see Section 2.3 for details). We assume the same convergence properties derived in Nadler et al. (2023b) and apply a cut on present-day subhalo virial mass of $M_{\mathrm{sub}}>300\times m_{\mathrm{part}}=1.2\times 10^{8}~{}{M_{\odot}}$ when analyzing individual simulations. We conservatively cut on $M_{\mathrm{sub}}/M_{\mathrm{host}}>1.4\times 10^{-4}$ when analyzing the entire suite; this threshold corresponds to $300\times m_{\mathrm{part}}$ for the lowest-mass Milky Way-est host.
@@ 2.2 Milky Way-est Host Halo Selection
    Host mass, $M_{\mathrm{host}}$, between $1\times 10^{12}$ and $1.8\times 10^{12}~{}{M_{\odot}}$ (corresponding to the $2\sigma$ range of MW host halo mass derived in Callingham et al. 2019);
@@
    Host concentration, $c_{\mathrm{host}}$, between 7 and 16 (corresponding to the $2\sigma$ range of MW host halo concentration derived in Callingham et al. 2019);
@@
    An LMC analog subhalo that satisfies:
@@
      Maximum circular velocity $V_{\text{max,sub}}>55~{}{\rm km\,s^{-1}}$ (van der Marel & Kallivayalil 2014; if multiple exist, we choose the subhalo with the highest peak maximum circular velocity, $V_{\text{peak,sub}}$);
@@
      Scale factor of most recent infall into the MW virial radius $a_{\mathrm{infall}}>0.86$ (lookback time within $1.3~{}{\rm Gyr}$ of $z=0$; Kallivayalil et al. 2013);
@@
    A GSE analog subhalo that merges with the MW host between $0.60.25<a_{\mathrm{disrupt}}<$ 0.6 with $M_{\mathrm{sub}}/M_{\mathrm{host}}>0.2$ when the GSE analog achieves its peak mass (e.g., Helmi 2020; Naidu et al. 2021).
These criteria are designed to bracket observationally inferred properties of the MW, LMC, and GSE halos. Starting with the host, our cuts encompass most MW halo mass and concentration estimates in the literature (e.g., see Wang et al. 2020 for a review). Note that we measure halo mass and concentration in the absence of baryons because we work with gravity-only simulations. Including baryons may affect the values of both quantities (e.g., Cautun et al. 2020), and it would be interesting to explore this effect with hydrodynamic resimulations of Milky Way-est hosts. Regardless, the MW halo properties from Callingham et al. (2019) that we use account for this effect.
Table 1: Percentage of MW-mass halos from c125-1024 that pass each Milky Way-est selection cut, corresponding to the criteria from Section 2.2. For each criterion, both the percentage of halos that pass only the previous cut and the remaining percentage of MW-mass halos are shown.
Criterion % of previous % of MW-mass
@@
Next, for the LMC, our $V_{\mathrm{max,sub}}$ cut is based on measurements of the LMC’s rotation curve (van der Marel & Kallivayalil, 2014). Although we do not place an additional cut on the LMC’s halo mass, we will show that our LMC analogs’ halo masses are broadly consistent with LMC dynamical mass measurements (see Section 3.2). [^2] Boylan-Kolchin et al. (2010) find that $\approx 10\%$ to $30\%$ of MW-mass halos host a subhalo with $M_{\mathrm{sub}}\gtrsim 10^{11}~{}{M_{\odot}}$. Some LMC analogs with $M_{\mathrm{sub}}<10^{11}~{}{M_{\odot}}$ satisfy our $V_{\mathrm{max,sub}}>55~{}{\rm km\,s^{-1}}$ criterion, which explains why a larger fraction of our hosts pass this cut than in Boylan-Kolchin et al. (2010); see Table 1. We impose a cut on $a_{\mathrm{infall}}$ so that our LMC analogs’ infall times are consistent with those derived from proper motion measurements (Kallivayalil et al., 2013). Note that most of the LMC analogs are on first infall; see Section 2.3 for details on the final sample. While the LMC’s current $50~{}{\rm kpc}$ distance from the MW is measured to within ±$\pm 1~{}{\rm kpc}$ (Pietrzyński et al., 2019), we include halos with LMC analogs within ±$\pm 20~{}{\rm kpc}$ of this distance at $z=0$ in order to obtain a statistically representative sample of LMC analogs. The LMC distance criterion is our most restrictive cut and yields LMC analogs that are broadly consistent with the LMC’s inferred orbit; we discuss this further in Section 3.2.
@@
We also explored including a Small Magellanic Cloud (SMC) analog. However, no halos that satisfy all cuts above have a clear SMC analog, i.e., a massive subhalo on a long-lived orbit around the LMC. In addition, we note that Milky Way-est halos are not explicitly selected to reside in environments representative of the Local Group or Local Volume; nonetheless, Milky Way-est hosts are generally less isolated than Symphony MW halos, which were selected with an explicit isolation cut (Nadler et al., 2023b). As a result, massive neighboring host halos reminiscent of M31 surround a handful of Milky Way-est systems. Studying the impact of Milky Way-est hosts’ environments on their accretion histories and subhalo populations is therefore an interesting area for future work.

### v2

@@ Abstract
We present Milky Way-est, a suite of 20 cosmological cold-dark-matter-only zoom-in simulations of Milky Way (MW)-like host halos. Milky Way-est hosts are selected such that they (i) are consistent with the MW’s measured halo mass and concentration, (ii) accrete a Large Magellanic Cloud (LMC)-like ($\approx 10^{11}~{}{M_{\odot}}$) subhalo within the last $2~{}{\rm Gyr}$ on a realistic orbit, placing them near $50~{}{\rm kpc}$ from the host center at $z\approx 0$, and (iii) undergo a $>$1:5 sub-to-host halo mass ratio merger with a Gaia–Sausage–Enceladus (GSE)-like system at early times ($30.67<z<$ 3). Hosts satisfying these LMC and GSE constraints constitute $<1\%$ of all halos in the MW’s mass range, and their total masses grow rapidly at late times due to LMC analog accretion. Compared to hosts of a similar final halo mass that are not selected to include LMC and GSE analogs, Milky Way-est hosts contain $22\%$ more subhalos with present-day virial masses above $10^{8}~{}{M_{\odot}}$ throughout the virial radius, on average. This enhancement reaches $\approx 80\%$ in the inner $100~{}\mathrm{kpc}$ and is largely, if not entirely, due to LMC-associated subhalos. These systems also induce spatial anisotropy in Milky Way-est subhalo populations, with $\approx 60\%$ of the total subhalo population within $100~{}{\rm kpc}$ found in the current direction of the LMC. Meanwhile, we find that GSE-associated subhalos do not significantly contribute to present-day Milky Way-est subhalo populations. These results provide context for our Galaxy’s dark matter structure and subhalo population and will help interpret a range of measurements that are currently only possible in the MW.
@@ 1 Introduction
The Milky Way (MW)’s dark matter distribution is a cornerstone of near-field cosmology. In recent years, astrometric measurements (Gardner et al., 2021), photometric surveys (Drlica-Wagner et al., 2020, 2022), and spectroscopic campaigns (Majewski et al., 2017; Naidu et al., 2020; Cooper et al., 2023) have qualitatively advanced our understanding of the events that contributed to the buildup of dark matter in the MW. These observations reveal that the MW’s most massive surviving and disrupted satellites—including the recently accreted Large Magellanic Cloud (LMC) system (van der Marel et al. 2002; Besla et al. 2007; Kallivayalil et al. 2013; Vasiliev 2023) and the ancient Gaia–Sausage–Enceladus (GSE) merger (Belokurov et al. 2018; Helmi et al. 2018)—play a crucial role in shaping its dark matter distribution and substructure (e.g., Helmi 2020; Conroy et al. 2021).
@@
In particular, the recently accreted LMC drives the MW’s dark matter distribution away from dynamical equilibrium, inducing wakes, deformation, and reflex motion that have recently been detected (e.g., Garavito-Camargo et al. 2019, 2021a; Petersen & Peñarrubia 2021; Conroy et al. 2021; Lilleengen et al. 2023). The LMC also brings its own substructure into the MW, reflected by the existence of kinematically associated satellite galaxies (e.g., Kallivayalil et al. 2018; Patel et al. 2020; Samuel et al. 2021) and spatial anisotropy in the distribution of MW satellite galaxies (e.g., Jethwa et al. 2016; Nadler et al. 2020). Meanwhile, the ancient GSE merger is a key event in the formation of the MW galaxy, as evidenced by the chemo-kinematic structure of the MW’s disk (e.g., Grand et al. 2020; Belokurov & Kravtsov 2022; Semenov et al. 2024; Dillamore et al. 2024). GSE also significantly impacts the dynamical structure of the MW stellar halo, which is dominated by stars on radial orbits deposited by this merger at early times (e.g., Fattahi et al. 2019; Dillamore et al. 2022; Han et al. 2022). Finally, the Sagittarius (Sgr) dwarf galaxy (Ibata et al. 1994) fell into the MW between the GSE and LMC and is tidally disrupting today (e.g., Vasiliev & Belokurov 2020). Sgr plays an important role in shaping the MW’s stellar halo, stellar streams population, and disk, and may have had an infall mass comparable to the LMC (e.g., Kruijssen et al. 2020).
@@
From a theoretical perspective, cosmological simulations provide detailed predictions for the growth and structure of MW-mass halos, including their subhalo populations (e.g., Springel et al. 2008; Diemand et al. 2008; Mao et al. 2015; Griffen et al. 2016; Sawala et al. 2016; Wetzel et al. 2016; Poole-McKenzie et al. 2020; Nadler et al. 2023b). Most studies that use simulations to interpret MW observations select hosts with masses and concentrations consistent with those inferred for the MW halo, regardless of merger history. In hydrodynamic contexts, selection often includes central galaxy properties (e.g., stellar mass and morphology). Other simulations require MW analogs to reside in environments similar to the Local Group or Local Volume (e.g., Garrison-Kimmel et al. 2014; Carlesi et al. 2016; Libeskind et al. 2020; Sawala et al. 2022).
@@
Motivated by the observational advances discussed above, recent simulations have focused on halos that satisfy additional constraints on the MW’s merger history. Considering the LMC, several authors have analyzed small numbers of MW-mass hosts with LMC analog subhalos using gravity-only (e.g., Erkal & Belokurov 2020; Sales et al. 2017) or hydrodynamic simulations (e.g., Barry et al. 2023; Smith-Orlik et al. 2023; Arora et al. 2023). Considering GSE and other early accretion events, Bignone et al. (2019) and Bose et al. (2020) study hosts in cosmological hydrodynamic simulations that include GSE analogs, Rey et al. (2023) present zoom-in simulations that are engineered to include GSE analogs with controlled properties, and García-Bethencourt et al. (2023) present a high-resolution hydrodynamical simulation that includes analogs of the GSE and the early Kraken and Sequoia mergers. Beyond individual constraints, Evans et al. (2020) analyze MW-like hosts with both LMC and GSE analogs using large-volume cosmological simulations, and Nadler et al. (2020) study zoom ins of two MW-mass hosts with analogs of both systems.
Despite this progress, no existing simulation suite is designed to study MW-mass hosts that accrete an LMC analog at late times and have an early accretion history consistent with that of the MW (e.g., by undergoing a major merger with a GSE analog at early times). Thus, it is timely to develop high-resolution cosmological simulations that incorporate such constraints at both early and late times.
@@
Here, we introduce Milky Way-est, a suite of $20$ cosmological cold-dark-matter-only zoom-in simulations focused on hosts that are consistent with the MW’s halo mass and concentration measurements and that accrete both LMC and GSE analogs. While we do not explicitly include an Sgr constraint in order to obtain a reasonably large host sample, we note that nearly half of our hosts accrete a massive subhalo between GSE disruption and LMC infall that merges by $z=0$. To benchmark our results, we compare Milky Way-est to the Symphony Milky Way-mass zoom-in simulation suite (hereafter Symphony MW; Mao et al. 2015; Nadler et al. 2023b). Symphony MW includes 45 MW-mass hosts, run with the same resolution and numerical settings as Milky Way-est, which are not selected to include LMC and GSE analogs. Comparing Milky Way-est and Symphony MW hosts can, therefore, help isolate the effects of our GSE and LMC constraints on hosts’ mass accretion histories (MAHs) and subhalo populations. We will show that Milky Way-est hosts’ formation times, subhalo mass functions (SHMFs), and subhalo radial and spatial distributions significantly differ from Symphony MW. Thus, it is important to model both the LMC and GSE to accurately interpret measurements of the MW’s dark matter distribution and substructure.
@@
Milky Way-est simulations resolve subhalos down to present-day masses of $\approx 10^{8}~{}{M_{\odot}}$, near the threshold of galaxy formation (Benitez-Llambay & Frenk, 2020; Nadler et al., 2020); detailed comparisons between Milky Way-est and Symphony MW will therefore help contextualize upcoming observations of the MW satellite population, which probe galaxy formation and dark matter physics (Nadler et al., 2024). More generally, Milky Way-est provides a realistic setting for testing empirical galaxy–halo connection models (e.g., Nadler et al. 2019; Wang et al. 2021), conducting zoom-in simulations of galaxy formation (e.g., Wetzel et al. 2016; Font et al. 2020; Applebaum et al. 2021), and exploring physics beyond the cold-dark-matter paradigm (e.g., Nadler et al. 2021b, a; Mau et al. 2022; Yang et al. 2023) in a setting that facilitates accurate predictions for near-field cosmology.
@@
This paper is organized as follows. Section 2 describes our simulations, including host halo selection criteria and the Milky Way-est sample. In Section 3, we present the properties of Milky Way-est hosts, LMC analogs, and GSE analogs. In Section 4, we study Milky Way-est host halo MAHs and subhalo populations compared to Symphony MW. We discuss the implications of our results, areas for future work, and caveats in Section 5. We conclude in Section 6. We list the names and properties of Milky Way-est systems (including their LMC and GSE analog halos) in Appendix A, and a convergence test using one higher-resolution Milky Way-est resimulation is presented in Appendix B.
@@ 2 Simulations
We begin by describing our simulations, including technical details (Section 2.1), host halo selection criteria (Section 2.2), the Milky Way-est sample that we resimulate using the zoom-in technique (Section 2.3), and the MW-mass sample, Symphony MW, that we use in comparative analyses with our suite (Section 2.4).
2.1 Simulation Setup
@@
Following Nadler et al. (2023b), all of our simulations use $h=0.7$, $\Omega_{\rm m}=0.286$, $\Omega_{\Lambda}=0.714$, $\sigma_{8}=0.82$, and $n_{s}=0.96$ (similar to WMAP 9; Hinshaw et al. 2013). Halo catalogs and merger trees are generated using Rockstar and Consistent-Trees (Behroozi et al., 2013a, b). We define “peak” as the highest value of a quantity along a halo’s main branch, and we measure concentration as $c_{\mathrm{vir}}\equiv R_{\mathrm{vir}}/R_{s}$, where $R_{s}$ is the Navarro–Frenk–White (Navarro et al. 1997) scale radius fit by Rockstar.
@@
Milky Way-est hosts are selected from Chinchilla c125-1024, a cosmological dark-matter-only simulation run with a box length of $125~{}{\rm Mpc}~{}h^{-1}$ and $1024^{3}$ particles (see Mao et al. 2015; Wang et al. 2021, for details). We select hosts based on several criteria, including host halo properties that are consistent with inferred properties of the MW’s dark matter halo and the existence of realistic LMC and GSE analogs, as described in Section 2.2. We generate zoom-in initial conditions with four refinement regions, yielding an equivalent of $8192$ particles per side in the most refined region using MUSIC (Hahn & Abel, 2011). The highest-resolution region corresponds to the Lagrangian volume of particles within 10 times the $z=0$ virial radius of each host ($R_{\mathrm{vir,host}}$) in c125-1024 (with the exception of Halo756, for which we use $7R_{\mathrm{vir,host}}$ to avoid an expensive resimulation). [^1] Virial quantities are calculated using the Bryan & Norman (1998) overdensity, which corresponds to $\Delta_{\mathrm{vir}}\approx 99$ times the critical density of the universe in our cosmology at $z=0$.
@@
We run zoom-in simulations using GADGET-2 (Springel, 2005), with numerical parameters matched to those for Symphony MW (see Mao et al. 2015; Nadler et al. 2023b for details). The dark matter particle mass in the highest-resolution region is $m_{\mathrm{part}}=4.0\times 10^{5}~{}{M_{\odot}}$ and the comoving Plummer-equivalent gravitational softening is $\epsilon=170~{}{\rm pc}~{}h^{-1}$. We also verify convergence using a resimulation of one Milky Way-est host with 8×8$\times$8 × smaller particle mass and 2×2$\times$2 × smaller softening (see Appendix B). We initialize each zoom in at $z=99$ and save 236 snapshots until $z=0$. At late times, the typical time spacing between snapshots is $25~{}{\rm Myr}$; as a result, we run some Milky Way-est zoom ins slightly past $z=0$ to ensure that our LMC analogs reach a distance of $\approx 50~{}{\rm kpc}$ (see Section 2.3 for details). We assume the same convergence properties derived in Nadler et al. (2023b) and apply a cut on a present-day subhalo virial mass of $M_{\mathrm{sub}}>300\times m_{\mathrm{part}}=1.2\times 10^{8}~{}{M_{\odot}}$ when analyzing individual simulations. We conservatively cut on $M_{\mathrm{sub}}/M_{\mathrm{host}}>1.4\times 10^{-4}$ when analyzing the entire suite, where $M_{\mathrm{host}}$ is defined as the total dark matter mass within $R_{\mathrm{vir,host}}$, including all subhalos. This threshold corresponds to $300\times m_{\mathrm{part}}$ for the lowest-mass Milky Way-est host.
@@ 2.2 Milky Way-est Host Halo Selection
    Host mass, $M_{\mathrm{host}}$, between $1\times 10^{12}$ and $1.8\times 10^{12}~{}{M_{\odot}}$ (corresponding to the $2\sigma$ range of MW host halo mass derived in Callingham et al. 2019).
@@
    Host concentration, $c_{\mathrm{host}}$, between 7 and 16 (corresponding to the $2\sigma$ range of MW host halo concentration derived in Callingham et al. 2019).
@@
    An LMC analog subhalo that satisfies the following:
@@
      Maximum circular velocity $V_{\text{max,sub}}>55~{}{\rm km\,s^{-1}}$ (van der Marel & Kallivayalil 2014; if multiple exist, we choose the subhalo with the highest peak maximum circular velocity, $V_{\text{peak,sub}}$).
@@
      Scale factor of most recent infall into the MW virial radius $a_{\mathrm{infall}}>0.86$ ($z_{\mathrm{infall}}<0.16$, corresponding to a lookback time within $2~{}{\rm Gyr}$ of $z=0$; Kallivayalil et al. 2013).
@@
    A GSE analog subhalo that merges with the MW host between $0.60.25<a_{\mathrm{disrupt}}<$ 0.6 (i.e., $30.67<z_{\mathrm{disrupt}}<$ 3 or between $6$ and $11.5~{}{\rm Gyr}$ ago) with $M_{\mathrm{sub}}/M_{\mathrm{host}}>0.2$ when the GSE analog achieves its peak mass (e.g., Helmi 2020; Naidu et al. 2021).
The selection criteria are designed to bracket observationally inferred properties of the MW, LMC, and GSE halos. Given these constraints, we find that $<1\%$ of MW-mass halos undergo both accretion events, adding to previous findings regarding the rarity of the LMC/Small Magellanic Cloud (SMC; e.g., Boylan-Kolchin et al. 2010; Busha et al. 2011; Liu et al. 2011).
Starting with the host, our cuts encompass most MW halo mass and concentration estimates in the literature (e.g., see Wang et al. 2020, for a review). Note that we measure halo mass and concentration in the absence of baryons because we work with gravity-only simulations. Including baryons may affect the values of both quantities (e.g., Cautun et al. 2020), and it would be interesting to explore this effect with hydrodynamic resimulations of Milky Way-est hosts. Regardless, the MW halo properties from Callingham et al. (2019) that we use account for this effect.
Table 1: Percentage of MW-mass halos from c125-1024 that pass each Milky Way-est selection cut, corresponding to the criteria from Section 2.2. For each criterion, the percentage of halos that pass only the previous cut and the remaining percentage of MW-mass halos are shown.
Criterion % of Previous % of MW Mass
@@
Next, for the LMC, our $V_{\mathrm{max,sub}}$ cut is based on measurements of the LMC’s rotation curve (van der Marel & Kallivayalil, 2014). Although we do not place an additional cut on the LMC’s halo mass, we will show that our LMC analogs’ halo masses are broadly consistent with LMC dynamical mass measurements (see Section 3.2). [^2] Boylan-Kolchin et al. (2010) find that $\approx 10\%$ to $30\%$ of MW-mass halos host a subhalo with $M_{\mathrm{sub}}\gtrsim 10^{11}~{}{M_{\odot}}$. Some LMC analogs with $M_{\mathrm{sub}}<10^{11}~{}{M_{\odot}}$ satisfy our $V_{\mathrm{max,sub}}>55~{}{\rm km\,s^{-1}}$ criterion, which explains why a larger fraction of our hosts pass this cut than in Boylan-Kolchin et al. (2010); see Table 1. We impose a cut on $a_{\mathrm{infall}}$ so that our LMC analogs’ infall times are consistent with those derived from proper-motion measurements (Kallivayalil et al., 2013). Most of the LMC analogs are on first infall; see Section 2.3 for details on the final sample. While the LMC’s current $50~{}{\rm kpc}$ distance from the MW is measured to within ±$\pm 1~{}{\rm kpc}$ (Pietrzyński et al., 2019), we include halos with LMC analogs within ±$\pm 20~{}{\rm kpc}$ of this distance at $z=0$ in order to obtain a statistically representative sample of LMC analogs. The LMC distance criterion is our most restrictive cut and yields LMC analogs that are broadly consistent with the LMC’s inferred orbit; we discuss this further in Section 3.2.